OpenStudy (danieke):

HELP: How do you find the solutions to a system of equations involving quadratic functions and other functions?

OpenStudy (anonymous):

what grade is this

OpenStudy (anonymous):

pemdas

OpenStudy (danieke):

12

OpenStudy (anonymous):

I am in 7th

OpenStudy (anonymous):

8th

OpenStudy (anonymous):

Example: Solve these two equations: •y = x2 - 5x + 7 •y = 2x + 1 Make both equations into "y=" format: They are both in "y=" format, so go straight to next step Set them equal to each other x2 - 5x + 7 = 2x + 1 Simplify into "= 0" format (like a standard Quadratic Equation) Subtract 2x from both sides: x2 - 7x + 7 = 1 Subtract 1 from both sides: x2 - 7x + 6 = 0

OpenStudy (anonymous):

Solve the Quadratic Equation! (The hardest part for me) You can read how to solve Quadratic Equations, but here we will factor the Quadratic Equation: Start with: x2 - 7x + 6 = 0 Rewrite -7x as -x-6x: x2 - x - 6x + 6 = 0 Then: x(x-1) - 6(x-1) = 0 Then: (x-1)(x-6) = 0 Which gives us the solutions x=1 and x=6

OpenStudy (anonymous):

do yall need help

OpenStudy (anonymous):

yeaa

OpenStudy (anonymous):

ok it is long

OpenStudy (danieke):

it is ?

OpenStudy (anonymous):

is this ur qusetion Algebra a system of equations Make both equations into "y =" format Set them equal to each other Simplify into "= 0" format (like a standard Quadratic Equation) Solve the Quadratic Equation! Use the linear equation to calculate matching "y" values, so we get (x,y) points as answers An example will help: Example: Solve these two equations: y = x2 - 5x + 7 y = 2x + 1 Make both equations into "y=" format: They are both in "y=" format, so go straight to next step Set them equal to each other x2 - 5x + 7 = 2x + 1 Simplify into "= 0" format (like a standard Quadratic Equation) Subtract 2x from both sides: x2 - 7x + 7 = 1 Subtract 1 from both sides: x2 - 7x + 6 = 0 Solve the Quadratic Equation! (The hardest part for me) You can read how to solve Quadratic Equations, but here we will factor the Quadratic Equation: Start with: x2 - 7x + 6 = 0 Rewrite -7x as -x-6x: x2 - x - 6x + 6 = 0 Then: x(x-1) - 6(x-1) = 0 Then: (x-1)(x-6) = 0 Which gives us the solutions x=1 and x=6

OpenStudy (anonymous):

was that it

OpenStudy (danieke):

yeah thank you so much guys ;)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

ur welcome

OpenStudy (danieke):

i have a next question tho

OpenStudy (anonymous):

tag mehhhh

OpenStudy (anonymous):

ok whats up